A firm has the following total costs, where Q is output and TC is total cost:...

  1. A firm has the following total costs, where Q is output and TC is total cost:

Q

TC

0

$ 100

1

110

2

130

3

160

4

200

5

250

6

310

7

380

8

460

9

550

10

650

11

760

  1. Say the firm is in a perfectly competitive market. If the current market (equilibrium) price is $ 70, at what output level will the firm as a profit maximizer produce at?

  1. Say the market price rises to $ 90. At what output level (as a perfect competitor) will this produce at?

  1. How much profit is the firm making at a price of $90? Based on this calculation, do you expect firms to enter or leave this market?

  1. Say instead this firm is a monopoly. If the firm maximizes profit at an output level where marginal revenue equals $ 60, what output level will this be?
Answer #1
Q MC = Change in TC
0 -
1 110 - 100 = 10
2 130 - 110 = 20
3 30
4 40
5 50
6 60
7 70
8 80
9 90
10 100
11 110

a) P = 70 = MC at Q = 7. So, perfectly competitive profit maximizer produces 7 units because it maximizes profit at the point where P = MC.

b) P = 90 = MC at Q = 9. So, perfectly competitive profit maximizer produces 9 units because it maximizes profit at the point where P = MC.

c) Profit = Total Revenue - TC = P*Q - TC = (90*9) - 550 = 810 - 550 = 260. As economic profit is positive, so firms will enter this market.

d) Monopolist maximizes profit where MR = MC. MR = MC = 60 at 6 units. So, monopolist will produce 6 units.